# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 2.3 McCann’s results for 2-body systems
Theorem 2.25 (vi) is slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vi)], which states “where $\rho$ is positive, it has as many as derivatives as the inverse of $A^{\prime}(\rho)$”. In McCann’s proof, he said “Because $V_{\rho}$ gains a derivative from $\rho$, smoothness of $\rho$ where positive follows from a bootstrap in (15)”, where “(15)” is (EL) in this paper. It is true that once $\rho\in W^{k,q}\left(\mathbb{R}^{3}\right)$ with some $q>3$ then we can obtain $V_{\rho}\in W^{k+1,\infty}\left(\mathbb{R}^{3}\right)\cap C^{k+1}\left(\mathbb{R}^{3}\right)$ by the similar arguments in Proposition A.11. But so far we only know $\rho\in C^{1}\left(\left\{{\rho>0}\right\}\right)\cap C_{c}^{0}\left(\mathbb{R}^{3}\right)$, one thing that is not clear is whether $\rho\in W^{1,q}\left(\{\rho>0\}\right)$, and hence whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$. The difficulty lies in the boundary behaviour: although $\rho$ is globally continuous and vanishes on $\partial\{\rho>0\}$, we do not know the regularity of $\partial\{\rho>0\}$ and we cannot rule out the possibility that $\nabla\rho$ blows up near the boundary. It means we might not be able to apply the result of trace operator (e.g., [25, Exercise 15.26]) to glue $\rho\cdot\mathbf{1}_{\{\rho>0\}}$ and $\rho\cdot\mathbf{1}_{\{\rho=0\}}$ (here $\mathbf{1}_{E}$ denotes the indicator function of $E$).
@@
If we do not know whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$, we might also not know whether $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Even if we assume, for example, $\phi\in C^{2}\left((0,\infty)\right)$, we know $\rho(x)=\phi\circ W_{\rho}(x)=\phi\circ\left(\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x\right)+V_{\rho}(x)+\lambda_{i}\right)$ on $\Omega_{i}$ due to Remark 2.26, but it might not be enough to show $\rho\in C^{2}(\{\rho>0\})$ if we do not know $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Therefore, it is not quite convincing if $\phi\in C^{r}((0,\infty))$ implies $\rho\in C^{r}(\{\rho>0\})$ if $r\geq 2$.
Remark 3.4.
Theorem 2.25 (vi) suggests $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ implies $\rho\in C^{1}(\{\rho>0\})$. But whether the converse is correct needs more discussion. For example, assume $A^{\prime\prime}(s)=0$ for some $s>0$, then the “derivative” of $\phi$ at point $t:=A^{\prime}(s)$ should be $\infty$, or more precisely, $\lim\limits_{h\rightarrow 0}\frac{\phi(t+h)-\phi(t)}{h}=\infty$. To show it, one can following the proof of inverse function theorem [38, Theorem 10.4.2] and notice $A^{\prime}$ and $\phi$ are strictly increasing.
Assume there is a point $x_{0}\in\{\rho>0\}$ such that $W_{\rho}(x_{0})=\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x_{0}\right)+V_{\rho}\left(x_{0}\right)+\lambda_{i}=t$, i.e. $\phi\circ W_{\rho}(x_{0})=\rho(x_{0})$ due to Remark 2.26, one may guess the gradient of $\rho$ does not exist at $x_{0}$, but could have problem if $\nabla W_{\rho}=\nabla\left(\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}\right)$ happens to be 0 at $x_{0}$. A trivial example which helps to describe such situation is $h(s)=g\circ f(s)$, where $g(t)=\left\{\begin{array}[]{ll}-\sqrt{-t},&t<0\\ \sqrt{t},&t\geq 0\end{array},f(s)=0\right$.. Then $h(s)=0$ has derivative at $s_{0}=0$, although the “derivative” of $g$ at $t_{0}=f\left(s_{0}\right)=$ 0 is $\infty$. Therefore, it is not obvious whether $\rho\in C^{1}(\{\rho>0\})$ can implies $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$.
@@
Fortunately, when we try to show $\rho$ satisfies (EP’), we do not necessarily need to deal with the gradient of $\rho$; we only need to consider the gradient of $P(\rho)$. As a result, (F4) can be relaxed to (F4’). Under the conditions of (F4’), we can still prove gradient of $P(\rho)$ exists and then $\rho$ satisfies (EP’), as shown in the proof of Part (vii) of Theorem 2.25 below.
Remark 3.5.
It’s important to note that it’s because $\rho$ is a $W^{\infty}$-local energy minimizer that we can extend (EL) to a neighborhood of $\{\rho>0\}$, and then obtain the global continuity Part (iii). It helps to exclude the case that minimizer (star) has non-zero pressure on the boundary. Moreover, this will be crucial in showing the existence of $\nabla P(\rho)$ on the boundary $\partial\{\rho>0\}$, and then Part (vii) in the following.
@@
$\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$ implies $\nabla V_{\rho}$ exists and is continuous in $\mathbb{R}^{3}$ thanks to Proposition A.11, so is $\nabla W_{\rho}$, where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}$ as in the proof of Part (iii-vi). In the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $W_{\rho}>0$. $\nabla W_{\rho}$ exists means $\nabla A^{\prime}(\rho)$ exists. Thanks to Lemma 3.1, we know
@@
The points remain to be checked is those on the boundary $\partial\{\rho=0\}=\partial\{\rho>0\}$. Since $\rho$ is continuous, we know $\rho=0$ on $\partial\{\rho=0\}$ and then similarly as above we have $\omega^{2}\rho(x)P_{12}(x)=-\rho(x)\nabla V_{\rho}(x)=0$. But to show (EP’) holds, we still need to check if $\nabla P(\rho(x))$ exists and is 0 on $\partial\{\rho=0\}$.
@@
Then, by Lemma 3.1 and Lagrange’s mean value theorem [35, 5.8 Theorem], $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$, such that
@@
Theorem 2.25 (vii) is also slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vii)], which states “If $P(\rho)$ is continuously differentiable on $0,\infty)$, then $\rho$ satisfies (EP’)”. Although there is only two-line proof of Part (vii) in McCann’s paper [31, Theorem 2.1]: “If $P(\rho)$ is continuously differentiable, then $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}$ and (vii) follows by taking the gradient of (15)”, where “(15)” is (EL) in this thesis, it is not quite trivial since we do not know whether $\nabla P(\rho)$ truly exists in $\mathbb{R}^{3}$ ((EL) holds only in a $\delta$-neighborhood). Moreover, even if $P$ is differentiable, the identity $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ for a scalar $s$ does not immediately imply $\nabla A^{\prime}(\rho)=\frac{\nabla P(\rho)}{\rho}$ for a function $\rho$. Indeed, if $\nabla\rho$ does not exist, expressions such as $\nabla A^{\prime}(\rho)=A^{\prime\prime}(\rho)\nabla\rho$ or $\frac{\nabla P(\rho)}{\rho}=\frac{P^{\prime}(\rho)\nabla\rho}{\rho}$ are not justified, let alone the possibility of canceling $\nabla\rho$ on both sides. We need more detailed discussions as above, under the assumption (F4’).
Remark 3.7.
When we show $\nabla P(\rho(x))$ exists on $\partial\{\rho>0\}$, one might want to follow the same arguments (chain rule) as in $\{\rho>0\}$, but there could be a problem that we might not have $\rho=\phi\circ W_{\rho}$ in a $\epsilon$-neighborhood of $\partial\{\rho>0\}$. However, it can still inspire us to consider the case $\rho\left(x_{n}\right)>0$ as above and follow the proof of chain rule.
Remark 3.8.
If (F4) holds, as we discuss in Remark 2.8, due to Part (iv) we know in the region $\{\rho>0\}$, $\nabla\rho$ exists, then $\nabla P(\rho)$ also exists and satisfies $\nabla P(\rho)=P^{\prime}(\rho)\nabla\rho=\rho A^{\prime\prime}(\rho)\nabla\rho=\rho\nabla A^{\prime}(\rho)$. This is what we show above, but the arguments here are easier, with this stronger assumption (F4), though such arguments may not be valid when we discuss the region $\{\rho=0\}$ or $\partial\{\rho>0\}$.
@@
In Remark 2.14 and Theorem 2.33, we mention that the support of constrained minimizer lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$ implies the constrained minimizer is local minimizer hence satisfies Theorem 2.25. In fact, if we revisits the proof of Theorem 2.25, those results (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) are also true without specifying it is a local minimizer. The crucial thing is deriving (EL) in a neighborhood of the support of constrained minimizer, which is true if the support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then we can verify (EP’) at the boundary of the support.
@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Let us first discuss the motivation of introducing $W^{\infty}$ metric: as McCann says in [31], when we discuss what local minima are, we should first specify the topology. The choice will be delicate. In order to have local minima, we hope the topology to be strong enough to preclude tunneling of mass (counterexample can be seen in [31, Example 3.6 and Remark 3.7]). On the other hand, to make the local minima physically meaningful, we hope the topology to be weak enough so that the evolution of physical flows is continuous (see Remark 2.30 or the text below). Thanks to some probability literatures, we can find a topology with these properties, that is, Wasserstein $L^{\infty}$ metric described e.g. by McCann [31], Ambrosio, Brué and Semola [2] or Givens and Shortt [18].
@@
$f_{\#}\tilde{\mu}(B):=\tilde{\mu}\left(f^{-1}(B)\right)\quad\forall B\in\mathcal{B}(Y)$
@@
It turns out $W^{\infty}(\rho,\kappa)$ is actually independent of the probability space $(S,\Sigma,v)$ we choose. Due to this fact, we can choose $S=[0,1]$ with Lebesgue measure, or $(S,\Sigma,v)=(X,\mathcal{B}(X),\rho)$. If we consider the latter case, we have
@@
The right side of the inequality can be viewed as a (generalized) Monge’s formulation of the optimal transport problem (see e.g. [2, Section 1.2 and Section 8.1]).
@@
Notice in [2, Section 8.1], the authors define another Wasserstein $L^{\infty}$ distance (with an abuse of notation denoted by $\widetilde{W}^{\infty}$) via transport plans and Kantorovich’s formulation of the optimal transport problem. In general, we also have $\widetilde{W}^{\infty}(\rho,\kappa)\leq\inf\left\{\|d(x,g(x))\|_{L^{\infty}(\rho)}\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\right\}$ [2, Section 2.2]. Moreover, when we consider Wasserstein $L^{p}$ distance given in [2], $p<\infty$, under additional assumption(s) if needed, we have $\widetilde{W}^{p}(\rho,\kappa)=(\inf\{\int_{X}d^{p}(x,g(x))d\rho(x)\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\})^{\frac{1}{p}}$ [2, Section 2.2 or Section 5.1]. However, the link between $\widetilde{W}^{\infty}$ and $W^{\infty}$ seems not clear. In the following, our arguments will be based on $W^{\infty}$’s definition.
Remark 4.5.
$\rho\in\mathcal{P}(X)$ may be represented in many ways such as the probability distribution. Consider $X=\mathbb{R}^{3}$ with $d(x,y)=|x-y|,\forall x,y\in\mathbb{R}^{3},\rho$ can be viewed as a law of a vector-valued random variable $f:S\rightarrow\mathbb{R}^{3}$ on the probability space $(S,\Sigma,v)$, with $f_{\#}v=\rho$. And, $W^{\infty}(\rho,\kappa)$ may now be defined as an infimum over all random variable representations of $\rho$ and $\kappa$ on a space $(S,\Sigma,v)$.
Remark 4.6.
Since the functions in ${R}\left(\mathbb{R}^{3}\right)$ have unit mass, we can view them as measures. That is, given $\rho\in{R}\left(\mathbb{R}^{3}\right)$, define $\widetilde{\rho}$ as the following: $\forall B\in\mathcal{B}\left(\mathbb{R}^{3}\right),\widetilde{\rho}(B):=\int_{B}\rho$. One can check in this definition, $\widetilde{\rho}\in\mathcal{P}\left(\mathbb{R}^{3}\right),\widetilde{\rho}$ is absolutely continuous with respect to $\mu$ (denoted by $\widetilde{\rho}\ll\mu$), where $\mu$ is Lebesgue measure, and the derivative (density) of $\widetilde{\rho}$ with respect to $\mu$ is $\rho$ [16, Section 1.6]. With an abuse of notation, we denote $\widetilde{\rho}$ by $\rho$ as well. Therefore, we can define the Wasserstein $L^{\infty}$ distance between two functions in ${R}\left(\mathbb{R}^{3}\right)$. Note that although $W^{\infty}(\rho,\kappa)$ may be infinite on ${R}\left(\mathbb{R}^{3}\right)$, it is finite whenever $\rho$ and $\kappa$ are of bounded support by Lemma 4.7 (i) below.
@@
We notice that the definition of Wasserstein distance can be generated to two functions in $t{R}\left(\mathbb{R}^{3}\right)=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=t\right\}$, with $t\neq 0$. Moreover, one can observe that given $\rho,\sigma$ in ${R}\left(\mathbb{R}^{3}\right),W^{\infty}(t\rho,t\sigma)=W^{\infty}(\rho,\sigma)$ (or $\widetilde{W}^{\infty}(t\rho,t\sigma)=\widetilde{W}^{\infty}(\rho,\sigma))$, which is different from Wasserstein $L^{p}$ distance.
@@
To prove ($vi$), the intuitive strategy is to “cut off” the function values of $\rho$ at the region where the value is relatively large, and then redistribute the value cut off to surrounding area where the values are relatively smaller, and then prove that the $W^{\infty}$ distance between the rearranged function and the original function is very small. To be precise, we carry out the following procedure. Without loss of generality we assume $t=1$, and we first divide the space $\mathbb{R}^{3}$ as $\mathbb{R}^{3}=\left(\cup_{n=1}^{\infty}C\left(x_{n}\right)\right)\cup D$. Here $\left\{C\left(x_{n}\right)\right\}$ is a family of (countably) many disjoint small open cubes, each with a diagonal length equal to $\frac{\epsilon}{2}$, volume equal to $V:=\left(\frac{\epsilon}{2\sqrt{3}}\right)^{3}$, centered at $x_{n}$. $D$ is the union of the remaining boundaries, $0\in D$, and $\mu(D)=0$. Since $\int\rho=1=\sum_{N=0}^{\infty}\int_{\{x\mid N\leq\rho(x)<N+1\}}\rho,\exists R>1$, such that $\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$. Let $\rho_{n}=\rho\cdot\mathbf{1}_{C\left(x_{n}\right)}$, where $\mathbf{1}_{C\left(x_{n}\right)}$ is indicator function of set $C\left(x_{n}\right)$, and define
$\sigma_{n}(x)=\left\{\begin{array}[]{ll}2R,&x\in\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x),&x\in\left\{x\mid R<\rho_{n}(x)<2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x)+\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)},&x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\\ 0,&x\in\mathbb{R}^{3}\backslash C\left(x_{n}\right)\end{array}\right$.
Since $\int_{\{x\mid\rho(x)>R\}}R<\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$, we know $\mu(\{x\mid\rho(x)>R\})<\frac{V}{4R}<\frac{V}{4}$, then
@@
$\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)\leq\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\rho_{n}\leq\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$
@@
$\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)}<\frac{1}{3}<\frac{3}{4}$

### v2

@@ 2.3 McCann’s results for 2-body systems
Theorem 2.22 (vi) is slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vi)], which states “where $\rho$ is positive, it has as many derivatives as the inverse of $A^{\prime}(\rho)$”. In McCann’s proof, he said “Because $V_{\rho}$ gains a derivative from $\rho$, smoothness of $\rho$ where positive follows from a bootstrap in (15)”, where “(15)” is (EL) in this paper. It is true that once $\rho\in W^{k,q}\left(\mathbb{R}^{3}\right)$ with some $q>3$ then we can obtain $V_{\rho}\in W^{k+1,\infty}\left(\mathbb{R}^{3}\right)\cap C^{k+1}\left(\mathbb{R}^{3}\right)$ by the similar arguments in Proposition B.3. But so far we only know $\rho\in C^{1}\left(\left\{{\rho>0}\right\}\right)\cap C_{c}^{0}\left(\mathbb{R}^{3}\right)$, one thing that is not clear is whether $\rho\in W^{1,q}\left(\{\rho>0\}\right)$, and hence whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$. The difficulty lies in the boundary behaviour: although $\rho$ is globally continuous and vanishes on $\partial\{\rho>0\}$, we do not know the regularity of $\partial\{\rho>0\}$ and we cannot rule out the possibility that $\nabla\rho$ blows up near the boundary. It means we might not be able to apply the result of trace operator (e.g., [25, Exercise 15.26]) to glue $\rho\cdot\mathbf{1}_{\{\rho>0\}}$ and $\rho\cdot\mathbf{1}_{\{\rho=0\}}$ (here $\mathbf{1}_{E}$ denotes the indicator function of $E$).
@@
If we do not know whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$, we might also not know whether $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Even if we assume, for example, $\phi\in C^{2}\left((0,\infty)\right)$, due to Remark 2.23, we know on $\Omega_{i}$,
$\rho(x)=\phi\circ W_{\rho}(x)=\phi\circ\left(\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x\right)+V_{\rho}(x)+\lambda_{i}\right)$
but it might not be enough to show $\rho\in C^{2}(\{\rho>0\})$ if we do not know $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Therefore, it is not quite convincing if $\phi\in C^{r}((0,\infty))$ implies $\rho\in C^{r}(\{\rho>0\})$ if $r\geq 2$.
Additionally, although Theorem 2.22 (vi) suggests $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ implies $\rho\in C^{1}(\{\rho>0\})$, whether the converse holds remains open.
@@
Note that when we try to show $\rho$ satisfies (EP’), we do not necessarily need to deal with the gradient of $\rho$; we only need to consider the gradient of $P(\rho)$. As a result, (F4) can be relaxed to (F4’). Under the conditions of (F4’), we can still prove the gradient of $P(\rho)$ exists and then $\rho$ satisfies (EP’), as shown in the proof of Part (vii) of Theorem 2.22 below.
@@
$\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$ implies $\nabla V_{\rho}$ exists and is continuous in $\mathbb{R}^{3}$ thanks to Proposition B.3, so is $\nabla W_{\rho}$, where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}$ as in the proof of Part (iii-vi). In the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $W_{\rho}>0$. $\nabla W_{\rho}$ exists means $\nabla A^{\prime}(\rho)$ exists. Thanks to Lemma 3.1, we know
@@
The points that remain to be checked are those on the boundary $\partial\{\rho=0\}=\partial\{\rho>0\}$. Since $\rho$ is continuous, we know $\rho=0$ on $\partial\{\rho=0\}$ and then similarly as above we have $\omega^{2}\rho(x)P_{12}(x)=-\rho(x)\nabla V_{\rho}(x)=0$. But to show (EP’) holds, we still need to check if $\nabla P(\rho(x))$ exists and is 0 on $\partial\{\rho=0\}$.
@@
Then, by Lemma 3.1 and Lagrange’s mean value theorem [35, Theorem 5.8], $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$, such that
@@
Theorem 2.22 (vii) is also slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vii)], which states “If $P(\rho)$ is continuously differentiable on $0,\infty)$, then $\rho$ satisfies (EP’)”. Although there is only two-line proof of Part (vii) in McCann’s paper [31, Theorem 2.1]: “If $P(\rho)$ is continuously differentiable, then $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}$ and (vii) follows by taking the gradient of (15)”, where “(15)” is (EL) in this paper, it is not quite trivial since we do not know whether $\nabla P(\rho)$ truly exists in $\mathbb{R}^{3}$ ((EL) holds only in a $\delta$-neighborhood). Moreover, even if $P$ is differentiable, the identity $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ for a scalar $s$ does not immediately imply $\nabla A^{\prime}(\rho)=\frac{\nabla P(\rho)}{\rho}$ for a function $\rho$. We need more detailed discussions as above, under the assumption (F4’).
@@
In Remark 2.13 and Theorem 2.30, we mention that the support of the constrained minimizer lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$ implies the constrained minimizer is a local minimizer, hence satisfies Theorem 2.22. In fact, if we revisit the proof of Theorem 2.22, those results (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) are also true without specifying it is a local minimizer. The crucial thing is deriving (EL) in a neighborhood of the support of the constrained minimizer, which is true if the support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then we can verify (EP’) at the boundary of the support.
@@ 4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
Let us first discuss the motivation of introducing $W^{\infty}$ metric: as McCann says in [31], when we discuss what local minima are, we should first specify the topology. The choice will be delicate. In order to have local minima, we hope the topology to be strong enough to preclude tunneling of mass (counterexample can be seen in [31, Example 3.6 and Remark 3.7]). On the other hand, to make the local minima physically meaningful, we hope the topology to be weak enough so that the evolution of physical flows is continuous (see Remark 2.27 or the text below). Thanks to some probability literature, we can find a topology with these properties, that is, the topology induced by Wasserstein $L^{\infty}$ metric. It is described, e.g., by McCann [31], Ambrosio, Brué and Semola [2] or Givens and Shortt [18].
@@
[…]
@@
It turns out $W^{\infty}(\rho,\kappa)$ is actually independent of the probability space $(S,\Sigma,v)$ we choose [31, Section 5]. Due to this fact, we can choose $S=[0,1]$ with Lebesgue measure, or $(S,\Sigma,v)=(X,\mathcal{B}(X),\rho)$. If we consider the latter case, we have
@@
The right side of the inequality can be viewed as a (generalized) Monge’s formulation of the optimal transport problem (see, e.g., [2, Section 1.2 and Section 8.1]).
@@
Notice in [2, Section 8.1], the authors define another Wasserstein $L^{\infty}$ distance (with an abuse of notation denoted by $\widetilde{W}^{\infty}$) via transport plans and Kantorovich’s formulation of the optimal transport problem. In general, we also have
$\widetilde{W}^{\infty}(\rho,\kappa)\leq\inf\left\{\|d(x,g(x))\|_{L^{\infty}(\rho)}\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\right\}$
see [2, Section 2.2]. Moreover, when we consider Wasserstein $L^{p}$ distance given in [2], $p<\infty$, under additional assumption(s) if needed, we have
$\widetilde{W}^{p}(\rho,\kappa)=\left(\inf\{\int_{X}d^{p}(x,g(x))d\rho(x)\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\}\right)^{\frac{1}{p}}$
see [2, Section 2.2 or Section 5.1]. However, the link between $\widetilde{W}^{\infty}$ and $W^{\infty}$ seems not clear. In the following, our arguments will be based on $W^{\infty}$’s definition.
Since the functions in ${R}\left(\mathbb{R}^{3}\right)$ have unit mass, we can view them as measures. That is, given $\rho\in{R}\left(\mathbb{R}^{3}\right)$, define $\widetilde{\rho}$ as the following: $\forall B\in\mathcal{B}\left(\mathbb{R}^{3}\right),\widetilde{\rho}(B):=\int_{B}\rho\,dx$. One can check in this definition, $\widetilde{\rho}\in\mathcal{P}\left(\mathbb{R}^{3}\right),\widetilde{\rho}$ is absolutely continuous with respect to $\mu$ (denoted by $\widetilde{\rho}\ll\mu$), where $\mu$ is Lebesgue measure, and the derivative (density) of $\widetilde{\rho}$ with respect to $\mu$ is $\rho$ [16, Section 1.6]. With an abuse of notation, we denote $\widetilde{\rho}$ by $\rho$ as well. Therefore, we can define the Wasserstein $L^{\infty}$ distance between two functions in ${R}\left(\mathbb{R}^{3}\right)$. Note that although $W^{\infty}(\rho,\kappa)$ may be infinite on ${R}\left(\mathbb{R}^{3}\right)$, it is finite whenever $\rho$ and $\kappa$ are of bounded support by Lemma 4.5 (i) below.
@@
We notice that the definition of Wasserstein distance can be generated to two functions in
$t{R}\left(\mathbb{R}^{3}\right)=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int_{\mathbb{R}^{3}}\rho\,dx=t\right\}$
Here $t\neq 0$. Moreover, one can observe that given $\rho,\sigma$ in ${R}\left(\mathbb{R}^{3}\right),W^{\infty}(t\rho,t\sigma)=W^{\infty}(\rho,\sigma)$ (or $\widetilde{W}^{\infty}(t\rho,t\sigma)=\widetilde{W}^{\infty}(\rho,\sigma))$, which is different from Wasserstein $L^{p}$ distance.
@@
To prove ($vi$), the intuitive strategy is to “cut off” the function values of $\rho$ at the region where the value is relatively large, and then redistribute the value cut off to surrounding area where the values are relatively smaller, and then prove that the $W^{\infty}$ distance between the rearranged function and the original function is very small. To be precise, we carry out the following procedure. Without loss of generality we assume $t=1$, and we first divide the space $\mathbb{R}^{3}$ as
$\mathbb{R}^{3}=\left(\cup_{n=1}^{\infty}C\left(x_{n}\right)\right)\cup D$
Here $\left\{C\left(x_{n}\right)\right\}$ is a family of (countably) many disjoint small open cubes, each with a diagonal length equal to $\frac{\epsilon}{2}$, volume equal to $V:=\left(\frac{\epsilon}{2\sqrt{3}}\right)^{3}$, centered at $x_{n}$. $D$ is the union of the remaining boundaries, $0\in D$, and $\mu(D)=0$. Since
$\int_{\mathbb{R}^{3}}\rho\,dx=1=\sum_{N=0}^{\infty}\int_{\{x\mid N\leq\rho(x)<N+1\}}\rho\,dx$
then $\exists R>1$, such that
$\int_{\{x\mid\rho(x)>R\}}\rho\,dx<\frac{V}{4}$
Let $\rho_{n}=\rho\cdot\mathbf{1}_{C\left(x_{n}\right)}$, where $\mathbf{1}_{C\left(x_{n}\right)}$ is indicator function of set $C\left(x_{n}\right)$, and define
$\sigma_{n}(x)=\left\{\begin{array}[]{ll}2R,&x\in\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x),&x\in\left\{x\mid R<\rho_{n}(x)<2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x)+\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)\,dx}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)},&x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\\ 0,&x\in\mathbb{R}^{3}\backslash C\left(x_{n}\right)\end{array}\right$.
Since $\int_{\{x\mid\rho(x)>R\}}R\,dx<\int_{\{x\mid\rho(x)>R\}}\rho\,dx<\frac{V}{4}$, we know $\mu(\{x\mid\rho(x)>R\})<\frac{V}{4R}<\frac{V}{4}$, then
@@
$\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)\,dx\leq\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\rho_{n}\,dx\leq\int_{\{x\mid\rho(x)>R\}}\rho\,dx<\frac{V}{4}$
@@
$\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)\,dx}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)}<\frac{1}{3}<\frac{3}{4}$
Therefore, if $x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)$, we get
$\sigma_{n}(x)\leq\rho_{n}(x)+\frac{3}{4}\leq R+\frac{3}{4}<2R$